40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 21 11:55:05 2015, Steve, Update, Optical Levers, 1103P noise measurement 1103P@qpdBS.jpgoplev1103P.pngoplevSum1103Pbs.png
    Reply  Tue Apr 21 15:03:27 2015, rana, Update, Optical Levers, 1103P noise measurement 
    Reply  Tue Apr 21 21:46:34 2015, rana, Update, Optical Levers, 1103P noise measurement oplev1103P.pngL1-ISI_SUS_ODC_FC496F_SPECTRUM-1113523216-86400.png
       Reply  Wed Apr 22 11:21:51 2015, Steve, Update, Optical Levers, BS & PRM oplevs are back to normal 
Message ID: 11231     Entry time: Tue Apr 21 15:03:27 2015     In reply to: 11230
Author: rana 
Type: Update 
Category: Optical Levers 
Subject: 1103P noise measurement 

It doesn't work with the lens in there, but it seems pretty close. Please leave it as is and I'll play with it after 5 today.

ELOG V3.1.3-